# Inferring Compositional 4D Scenes without Ever Seeing One

### v1

@@ Abstract
Scenes in the real world are often composed of several static and dynamic objects. Capturing their 4-dimensional structures, composition and spatio-temporal configuration in-the-wild, though extremely interesting, is equally hard. Therefore, existing works often focus on one object at a time, while relying on some category-specific parametric shape model for dynamic objects. This can lead to inconsistent scene configurations, in addition to being limited to the modeled object categories. We propose COM4D (Compositional 4D), a method that consistently and jointly predicts the structure and spatio-temporal configuration of 4D/3D objects using only static multi-object or dynamic single object supervision. We achieve this by a carefully designed training of spatial and temporal attentions on 2D video input. The training is disentangled into learning from object compositions on the one hand, and single object dynamics throughout the video on the other, thus completely avoiding reliance on 4D compositional training data. At inference time, our proposed attention mixing mechanism combines these independently learned attentions, without requiring any 4D composition examples. By alternating between spatial and temporal reasoning, COM4D reconstructs complete and persistent 4D scenes with multiple interacting objects directly from monocular videos. Furthermore, COM4D provides state-of-the-art results in existing separate problems of 4D object and composed 3D reconstruction despite being purely data-driven. Code will be released at https://github.com/insait-institute/COM4D.
@@ 4 Experiments
Geometric fidelity is measured using the Chamfer Distance (CD) [3, 5] and F-Score [77] (at a 0.1 threshold), where lower CD and higher F-Score are better. For 3D scenes, we concatenate all parts into a single mesh before evaluation, whereas for 4D sequences, we compute the metrics per-frame and then average. To assess structural quality, we use two Intersection-over-Union (IoU) metrics based on a consistent $64^{3}$ voxelization. In 3D, we measure part independence via a pairwise IoU, where lower scores indicate better separation between parts. In 4D, we evaluate accuracy with a per-frame IoU against the ground truth, where higher scores are better. Note that for Gaussian-based baselines (L4GM, GVFD), we convert their outputs to point clouds for CD and F-Score evaluation, but omit the IoU metric due to lack of reliable watertight meshes. All reported scores are the mean values over the test set.
@@ 4.4 Ablation Study
We report quantitative results in two forms. Because no compositional 4D dataset with object-level decomposition exists, our main evaluation relies on user studies in Inferring Compositional 4D Scenes without Ever Seeing One.B, where our method is preferred over a standard generative baseline (without Attention Mixing) by a factor of 12 (87% vs. 6.9%). We also evaluate on CMU Panoptic [32, 33] point clouds by registering only the first frame reconstruction, where Attention Mixing reduces the average CD from 35.91 cm to 7.42 cm. This demonstrates that COM4D can accurately capture multiple dynamic shape evolutions in long sequences >90>90 frames, with surprisingly small drifts. Fig. 4 visualizes the reconstructions of ours and the baseline. Further details appear in the supplementary material.
@@ Acknowledgements.
This research was partially funded by the Ministry of Education and Science of Bulgaria (support for INSAIT, part of the Bulgarian National Roadmap for Research Infrastructure). The research was also partly funded by INSAIT-VIVO project on 3D scene understanding.
@@ C Evaluation on CMU Panoptic Dataset
D Additional Qualitative Results on Compositional 4D
@@
E Additional Qualitative Results on 4D Object Reconstruction
@@
F Additional Qualitative Results on 3D Scene Generation

### v2

@@ Abstract
Scenes in the real world are often composed of several static and dynamic objects. Capturing their 4-dimensional structures, composition and spatio-temporal configuration in-the-wild, though extremely interesting, is equally hard. Therefore, existing works often focus on one object at a time, while relying on some category-specific parametric shape model for dynamic objects. This can lead to inconsistent scene configurations, in addition to being limited to the modeled object categories. We propose COM4D (Compositional 4D), a method that consistently and jointly predicts the structure and spatio-temporal configuration of 4D/3D objects using only static multi-object or dynamic single object supervision. We achieve this by a carefully designed training of spatial and temporal attentions on 2D video input. The training is disentangled into learning from object compositions on the one hand, and single object dynamics throughout the video on the other, thus completely avoiding reliance on 4D compositional training data. At inference time, our proposed attention mixing mechanism combines these independently learned attentions, without requiring any 4D composition examples. By alternating between spatial and temporal reasoning, COM4D reconstructs complete and persistent 4D scenes with multiple interacting objects directly from monocular videos. Furthermore, COM4D provides state-of-the-art results in existing separate problems of 4D object and composed 3D reconstruction despite being purely data-driven. Code is available at https://github.com/insait-institute/COM4D.
@@ 4 Experiments
Geometric fidelity is measured using the Chamfer Distance (CD) [3, 5] and F-Score [77] (at a 0.1 threshold), where lower CD and higher F-Score are better. For 3D scenes, we concatenate all parts into a single mesh before evaluation, whereas for 4D sequences, we compute the metrics per-frame and then average. To assess structural quality, we use two Intersection-over-Union (IoU) metrics based on a consistent $64^{3}$ voxelization. In 3D, we measure part independence via a pairwise IoU, where lower scores indicate better separation between parts. In 4D, we evaluate accuracy with a per-frame IoU against the ground truth, where higher scores are better. Note that for Gaussian-based baselines (L4GM, GVFD), we convert their outputs to point clouds for CD and F-Score evaluation, but omit the IoU metric for lack of reliable watertight meshes. All reported scores are the mean values over the test set.
@@ 4.4 Ablation Study
We report quantitative results in two forms. Because no compositional 4D dataset with object-level decomposition exists, our main evaluation relies on user studies in Inferring Compositional 4D Scenes without Ever Seeing One.B, where our method is preferred over a standard generative baseline (without Attention Mixing) by a factor of 12 (87% vs. 6.9%). We also evaluate on CMU Panoptic [32, 33] point clouds by registering only the first frame reconstruction, where Attention Mixing reduces the average CD from 35.91 cm to 7.42 cm. This demonstrates that COM4D can accurately capture multiple dynamic shape evolutions in long sequences >90>90 frames, with surprisingly small drifts. Fig. 4 visualizes the reconstructions of ours and the baseline. Further details and additional quantitive analysis on synthetic compositional data appear in the supplementary material.
@@ Acknowledgements.
This research was partially funded by the Ministry of Education and Science of Bulgaria (support for INSAIT, part of the Bulgarian National Roadmap for Research Infrastructure). The research was also partly funded by INSAIT-VIVO project on 3D scene understanding and editing.
@@ C Evaluation on CMU Panoptic Dataset
D Scalability with Respect to Object Count
COM4D supports denoising with up to 8 parts, and up to 16 parts with additional finetuning. In the static dataset, scenes contain on average 6.19 objects, with the most common configuration being 5 objects (16.67%).
We observe that performance is moderately sensitive to both the number of dynamic objects and the complexity of the static scene. As the number of interacting objects increases, the compositional reasoning task becomes more challenging, leading to gradual degradation in reconstruction quality.
E Additional Qualitative Results on Compositional 4D
@@
F Additional Qualitative Results with Moving Camera
COM4D assumes a fixed camera when reconstructing compositional scenes containing both static and dynamic objects, as camera motion is inherently entangled with static scene geometry. In such settings, disentangling camera motion from scene structure is fundamentally ambiguous without additional constraints or supervision.
That said, COM4D is not inherently restricted to static-camera scenarios. In cases where the scene consists only of dynamic objects, the model can still recover consistent relative 4D structure and motion over time. This is because COM4D relies on learning relative spatial relationships and temporal coherence, rather than absolute camera-referenced geometry.
Input
Refer to caption Refer to caption Refer to caption Refer to caption
Reconstr.
Refer to caption Refer to caption Refer to caption Refer to caption
Novel View
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 13: Qualitative results for a sample with dynamic camera.
G Physical Plausibility and Object Interactions
COM4D is a fully data-driven framework and does not explicitly enforce physical constraints such as contact or collision avoidance. Instead, interactions between multiple objects are modeled implicitly through attention mechanisms in the latent space.
To assess whether this implicit modeling yields physically plausible reconstructions, we evaluate mesh interpenetration across 8 video sequences by measuring the intersection-over-union (IoU) between reconstructed object meshes. We observe a low average overlap of $\text{IoU}=0.0096$, indicating minimal interpenetration between interacting objects.
For reference, we report an IoU of $0.0018$ on 3D-FRONT, where ground-truth scenes exhibit little to no physical contact between objects (see Tab. 2). The small gap between these values suggests that COM4D maintains a comparable level of physical plausibility despite not explicitly modeling physical constraints.
Figure 14: Qualitative results on interaction scenarios. Top row shows input frames, and bottom shows 4D reconstructions.
H Comparison with Optimization-Based Baselines
We compare COM4D against the recent optimization-based method Shape of Motion (SOM) [80]. SOM struggles to recover complete geometry under occlusions, often producing incomplete or noisy reconstructions (see Fig. 15).
Quantitatively, SOM achieves a Chamfer Distance (CD) of 11.30 cm, whereas COM4D achieves 7.42 cm on [32, 33]. This gap highlights the advantage of our generative formulation, which leverages learned spatial and temporal priors to produce more complete and coherent reconstructions in challenging compositional settings.
                       Input
Refer to caption Refer to caption Refer to caption Refer to caption
                  SOM [80]
Refer to caption Refer to caption Refer to caption Refer to caption
                   Ours
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 15: Qualitative comparison across methods.
I Evaluation Using a Synthetic Dataset
We additionally evaluate our method on a synthetic 4D dataset since a suitable real-captured 4D scene benchmark is unavailable. We first generate candidate synthetic scenes by composing one or more dynamic humanoid motion sequences with several static objects. Each dynamic object is selected from a motion sequence, sub-sampled with a fixed frame stride, and placed into a shared normalized world coordinate system. Static assets are then scaled, rotated around the vertical axis, and translated so that they fit around the dynamic trajectories without excessive overlap. For every time step, we export a combined GLB that contains the static geometry and the current pose of each dynamic object, and we also save metadata describing the source assets, transforms, per-frame bounds, and global centering applied to the scene.
For visualization, each frame is rendered from a fixed orthographic camera with consistent framing across the entire sequence. The renderer keeps the scene centered and uses the same camera and lighting conventions for all frames within a sample, which makes frame-to-frame motion directly comparable.
For quantitative evaluation, we compare the generated scene-level mesh against the static part of the reference scene and compute dynamic scores per moving object and per frame. For static geometry, we compute one CD/F-score pair per scene and report the arithmetic mean over scenes. For dynamic content, we first average all valid dynamic per-frame scores within each sample, and then average these per-sample dynamic means across the whole set. This prevents samples with more moving objects or more valid frames from dominating the final measure.
Component CD $\downarrow$ F-score $\uparrow$
Static scene geometry 0.293 0.418
Dynamic objects 0.284 0.434
Table 4: Quantitative results on the synthetic 4D evaluation set. Lower CD is better; higher F-score is better.
These synthetic-scene scores are worse than the corresponding single-object 4D and static compositional 3D scene results. The main reason is that, although the reconstructed objects themselves and their coarse arrangements are often visually consistent with the input, the relative positioning of objects is still somewhat inaccurate. Small errors in placement and pairwise spatial relationships accumulate at the full-scene level and are directly penalized by both CD and F-score, leading to lower quantitative performance than in simpler single-object or static-only settings.
J Additional Qualitative Results on Synthetic Compositional 4D Dataset
In Fig. 16, Fig. 17, Fig. 18 we show additional qualitative results on our synthetic compositional 4D dataset.
Input Refer to caption Refer to caption Refer to caption Refer to caption
COM4D Refer to caption Refer to caption Refer to caption Refer to caption
Input Refer to caption Refer to caption Refer to caption Refer to caption
COM4D Refer to caption Refer to caption Refer to caption Refer to caption
Input Refer to caption Refer to caption Refer to caption Refer to caption
COM4D Refer to caption Refer to caption Refer to caption Refer to caption
Input Refer to caption Refer to caption Refer to caption Refer to caption
COM4D Refer to caption Refer to caption Refer to caption Refer to caption
Figure 16: Additional qualitative results on the synthetic compositional 4D dataset.
Input Refer to caption Refer to caption Refer to caption Refer to caption
COM4D Refer to caption Refer to caption Refer to caption Refer to caption
Input Refer to caption Refer to caption Refer to caption Refer to caption
COM4D Refer to caption Refer to caption Refer to caption Refer to caption
Input Refer to caption Refer to caption Refer to caption Refer to caption
COM4D Refer to caption Refer to caption Refer to caption Refer to caption
Input Refer to caption Refer to caption Refer to caption Refer to caption
COM4D Refer to caption Refer to caption Refer to caption Refer to caption
Figure 17: Additional qualitative results on the synthetic compositional 4D dataset.
Input Refer to caption Refer to caption Refer to caption Refer to caption
COM4D Refer to caption Refer to caption Refer to caption Refer to caption
Input Refer to caption Refer to caption Refer to caption Refer to caption
COM4D Refer to caption Refer to caption Refer to caption Refer to caption
Input Refer to caption Refer to caption Refer to caption Refer to caption
COM4D Refer to caption Refer to caption Refer to caption Refer to caption
Input Refer to caption Refer to caption Refer to caption Refer to caption
COM4D Refer to caption Refer to caption Refer to caption Refer to caption
Figure 18: Additional qualitative results on the synthetic compositional 4D dataset.
K Additional Qualitative Results on 4D Object Reconstruction
@@ I Evaluation Using a Synthetic Dataset
L Additional Qualitative Results on 3D Scene Generation